OpenStudy (anonymous):

Kevin and Conner tracked their exercise last week. Kevin ran 8 times around the school track, 6 times around his neighborhood, and 7 miles on the treadmill. Conner ran 4 times around the school track and 19 miles on the treadmill. Write and simplify an expression for the total combined distance that Kevin and Conner ran. Show your work. Use t for the distance in miles around their school track and n for the distance in miles around Kevin’s neighborhood.

OpenStudy (anonymous):

8t+6n+7=4t+19 8t-4t+6n=19-7 4t+6n=12

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

can you answer this one